A truck travels for 4 hours at an average speed of 42 miles per hour. How much longer must the truck travel at an average speed of 55 miles per hour so that the average speed for the total trip is 50 miles per hour.
Answer by drk(1908)

This is an RTD problem. We can create the following table based on given information.
truck . . . . . .rate . . . . . .time . . . . . distance
truck . . . . . . 42 . . . . . . . .4 . . . . . . . .168
truck . . . . . . 55 . . . . . . . .x . . . . . . . .55x
totals. . . . . . 50. . . . . . . .4+x. . . . . . . 200 + 50x
looking at column three we get
168 + 55x = 200 + 50x
subtract 168 to get
55x = 32 + 50x
subtract 50x to get
5x = 32
divide to get
x = 6.4